OpenStudy (anonymous):

Put the equation below in the form ax2 + bx + c = 0. Enter exponents using the caret ( ^ ). For example, you would enter 4x2 as 4x^2. (x + 3)2 + 4x = 0

OpenStudy (anonymous):

@jim_thompson5910 can you help me please?

jimthompson5910 (jim_thompson5910):

(a+b)^2 = (a+b)(a+b) (a+b)^2 = a(a+b)+b(a+b) (a+b)^2 = a^2+ab+ab+b^2 (a+b)^2 = a^2+2ab+b^2

jimthompson5910 (jim_thompson5910):

So for example (not this problem, but something similar) If a = x and b = 7, then (a+b)^2 = a^2+2ab+b^2 (x+7)^2 = x^2+2x*7+7^2 (x+7)^2 = x^2+14x+49

OpenStudy (anonymous):

x^2+6x+9

jimthompson5910 (jim_thompson5910):

good, so (x + 3)^2 = x^2+6x+9 Now add 4x to both sides and tell me what you get

OpenStudy (anonymous):

4x^3+10x^2+15x+0 i feel like i did that wrong.

jimthompson5910 (jim_thompson5910):

6x+4x = ???

OpenStudy (anonymous):

10x

jimthompson5910 (jim_thompson5910):

So (x + 3)^2 = x^2+6x+9 (x + 3)^2 + 4x = x^2+6x+9 + 4x ... Add 4x to both sides. (x + 3)^2 + 4x = x^2 + (6x+4x) + 9 (x + 3)^2 + 4x = x^2 + (10x) + 9 (x + 3)^2 + 4x = x^2 + 10x + 9 Making sense?

OpenStudy (anonymous):

x^2+9+4x=2x^2+10x+9 x^2+5x=0?

jimthompson5910 (jim_thompson5910):

how are you getting that?

jimthompson5910 (jim_thompson5910):

oh is this a whole new problem???

OpenStudy (anonymous):

im soo confused

jimthompson5910 (jim_thompson5910):

did you just post a whole new problem?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

oh, then i have no idea how you got that

OpenStudy (anonymous):

i dont either i dont know what im doing.

jimthompson5910 (jim_thompson5910):

do you see how I got (x + 3)^2 + 4x = x^2 + 10x + 9 ???

OpenStudy (anonymous):

no.

jimthompson5910 (jim_thompson5910):

(step 1) (x + 3)^2 = x^2+6x+9 (step 2) (x + 3)^2 + 4x = x^2+6x+9 + 4x ... Add 4x to both sides. (step 3) (x + 3)^2 + 4x = x^2 + (6x+4x) + 9 (step 4) (x + 3)^2 + 4x = x^2 + (10x) + 9 (step 5) (x + 3)^2 + 4x = x^2 + 10x + 9 Tell me which step you're stuck at